# On the Robust Approximation of ASR Metrics

### v1

@@ 1 Introduction
Automatic Speech Recognition (ASR) models have made significant advancements in recent years, achieving near-human performance on several standard evaluation benchmarks (Radford et al., 2022; Seamless Communication et al., 2023; Communication et al., 2023; Harper et al., , inter alia). These models are typically evaluated using metrics like Word Error Rate (WER) and Character Error Rate (CER) Likhomanenko et al. (2020), which are essential for assessing model performance.
@@
  • We evaluate over 40 ASR models across 14 diverse evaluation setups, including both standard benchmarks and domain-specific, unseen conditions followed by training regression models to approximate ASR metrics.
  • We compare our approach with the most recent work on approximating ASR metrics and demonstrate over a 100% improvement against the strong baseline.
@@
Reproducibility. We are committed to making all code, data, and logs available upon acceptance.
@@ 2 Related Work
Automatic speech recognition (ASR) has seen remarkable progress in recent years, driven by advances in deep learning and the availability of extensive training datasets Radford et al. (2022); Communication et al. (2023). Transformer Vaswani et al. (2023) based models, in particular, have significantly contributed to these developments by effectively capturing long-range dependencies and contextual nuances in speech, achieving state-of-the-art (SOTA) performance across diverse benchmarks Kheddar et al. (2024); Dhanjal and Singh (2024); Zimerman and Wolf (2023). While traditional evaluation metrics like Word Error Rate (WER) and Character Error Rate (CER) are de-facto evaluation metrics in benchmarking ASR systems Lin et al. (2021); Park et al. (2024), scenarios where ground truth transcriptions are unavailable have caught interest in reference-free ASR evaluation methods Karbasi and Kolossa (2022); Wang et al. (2024); Kuhn et al. (2024).
Reference-free ASR evaluation methods aim to estimate ASR performance without requiring ground truth transcriptions Ospanov et al. (2024). Earlier approaches rely on heuristic features or metadata such as speaker demographics, background noise, and linguistic characteristics Litman et al. (2000); Yoon et al. (2010), limiting their applicability across varied contexts. However, recent advancements focus on deep learning-based frameworks, such as convolutional neural networks (CNNs) Elloumi et al. (2018) and contrastive learning methhods Yuksel et al. (2023a), to predict ASR quality directly from encoded speech and text. For instance, methods like NoRefER Yuksel et al. (2023b) employ Siamese architectures fine-tuned on ASR hypotheses, achieving high correlation with traditional metrics and improving WER by optimizing hypothesis ensembling Park et al. (2024).
@@ 3.4 Evaluation
We evaluate the regression model’s performance across four setups, including IID and OOD data and different model configurations. Specifically, we train our regression model on one ASR system (source) on one dataset and evaluate it on both IID and OOD data for the source and target models.
@@ 4.1 Datasets
To evaluate the robustness and generalizability of our ASR metrics approximation tool, we use datasets sourced from multiple distributions, divided into two types: Standard Benchmark and Wild Challenge datasets. Below we describe the datasets and provide additional details in Appendix A.2 Table 3.
@@ 4.2 Models
Encoder-Decoder Models. We include multiple encoder-decoder families of models capable of performing ASR tasks in a zero-shot setting. More specifically, we include whisper Radford et al. (2023) and distil-whisper Gandhi et al. (2023) models which perform really well across diverse testing settings. We also include seamless  Communication et al. (2023); Seamless Communication et al. (2023); Barrault et al. (2025), SpeechT5 Ao et al. (2022) which are unified encoder-decoder framework for tasks such as ASR, speech synthesis, translation, and voice conversion. MMS Pratap et al. (2023) supports hundreds of languages and excels in resource-constrained scenarios. Moonshine (2) Jeffries et al. (2024), a lightweight and efficient model, is designed for edge deployments with strong performance.
NeMo-ASR Models. We use multiple models from the NeMo-ASR Gulati et al. (2020); Variani et al. (2020); Noroozi et al. (2024); Tang et al. (2023); Harper et al. toolkit by NVIDIA. We include models such as Canary and Parakeet, which use highly efficient speech encoders like Fast-Conformer Rekesh et al. (2023). In addition to that, we use models based on various encoders, and decoders (CTC, RNN-T, TDT, Conformer-CTC Guo et al. (2021). In our work, we evaluate 11 models from the NeMo-ASR toolkit.
Encoder-Only and Decoder-Only Models. We include self-supervised encoder-only models and their derivatives, as well as decoder-only models like SpeechLLM. Specifically, we use Wav2Vec2 Schneider et al. (2019); Baevski et al. (2020), HuBERT Hsu et al. (2021), and Data2Vec Baevski et al. (2022). Additionally, we include speech language models like SpeechLLM Rajaa and Tushar , which combines speech embeddings with language models to predict metadata such as speaker attributes, emotions, and accents, offering robust multimodal capabilities.
@@ 4.3 Experimental Setup
For multimodal embeddings, we use SONAR Duquenne et al. (2023), a 1024-dimensional sentence-level multilingual embedding model. Specifically, we utilize text_sonar_basic_encoder for text encoding and speech_sonar_basic_encoder for speech encoding. These encoders provide unified representations.
@@ 5 Results
We conduct experiments using two dataset categories: standard benchmarks and in-the-wild, as described in Section 4.1. For each ASR model, a leave-one-out strategy is used, training the regression model on 9 benchmark datasets and testing it on the remaining benchmark dataset and all four in-the-wild datasets to ensure comprehensive evaluation exclusively on out-of-domain data. Additionally, in-domain testing is included in ablation studies, as detailed in Section 5.3. The regression model is trained to predict absolute error counts (word and character levels), which are normalized by the reference length to compute approximate error rates ($aWER$ and $aCER$). We also train regression models to directly predict WER and CER.
@@ 5.1 Evaluation on In-the-Wild Datasets
The wild datasets provide a realistic testbed for evaluating the regression model’s ability to approximate error rates under real-world conditions. The results are presented in Table 1. High-performing models, like canary-1b, demonstrate strong agreement between predicted and actual error rates. For example, on VP_Accented, canary-1b achieves a WER of $23.2\%$ and an $aWER$ of $12.1\%$, with a minimal difference of $1.1\%$. On Primock57, a clinical consultation dataset, the model shows robustness with a WER of $16.2\%$ and an $aWER$ of $13.4\%$, highlighting its effective generalization across diverse and domain-specific contexts.
For models like data2vec-audio-large-960h our approximation is pretty close to actual error rates with difference consistently under $2\%$ on various datasets. For example, on LibriSpeech-test-noise, the model’s actual WER is $7.2\%$ while the approximated $aWER$ is $8.6\%$, showcasing its reliability in noisy conditions. Even on acoustically complex datasets like ATCOsim, where the WER is $44.0\%$ and the $aWER$ is $51.1\%$, the model exhibits a reasonable alignment between approximated and actual error rates.
@@ 5.2 Evaluation on Benchmark Datasets
We summarize results on 10 standard benchmark datasets in Appendix A.5 Tables 8 and 9. Each table reports actual WER/CER alongside the approximated WER/CER (denoted by aWER/aCER).
Figure 2: Actural and approximated WER for four models across standard benchmark.
@@ 5.3 Ablation
We conduct ablation experiments to evaluate the robustness of the approximation model and the contributions of its individual components. Using the evaluation setup outlined in Section 3.4, we select data2vec-audio-base-960h as the source model ($S$) and wav2vec2-base-960h as the target model ($T$). The results are summarized in Table 5.3, where IID results correspond to Case-I 3.4, and $D$, $M$, and $D+M$ under OOD represent Case II 3.4, Case-III 3.4, and Case-IV 3.4, respectively. The reference model’s $r$ value represents the average WER across all datasets. We include reference models with varying $r$ values, such as whisper-large-v3 ($r=17.8$), whisper-medium.en ($r=20.1$), whisper-tiny ($r=33.4$), and mms-1b-fl102 ($r=51.0$).
@@ D M D + M
The quality of references, quantified by the $r$-value, also plays a critical role. For example, in IID conditions, the MAE increases from $1.31$ for $r=17.8$ to $2.03$ for $r=51.0$. A similar trend is observed in OOD $D+M$, where the MAE rises from $1.40$ ($r=17.8$) to $2.09$ ($r=51.0$). The absence of similarity (w/o S) combined with low-quality proxies further degrades performance, underscoring the importance of both high-quality references and similarity measures. These trends are similarly observed for character-level error count approximation, as detailed in Appendix Table 6.
@@ 6 Conclusion
We present a framework for approximating ASR metrics, demonstrating its effectiveness in generalizing to unseen, in-the-wild, and challenging conditions. Our results show that the model performs well with absolute error counts, consistently outperforming strong baseline, with error rates remaining relatively low. We show that our proposed method achieves consistent performance across 40 ASR models and 14 evaluation setups, including both standard benchmarks and domain-specific conditions. The trained regression model can be efficiently used to approximate ASR metrics, particularly in data-constrained environments, such as critical domains with limited labeled data. In summary, our work bridges the gap between theoretical advancements and real-world applications, paving the way for more reliable and scalable ASR systems. While in this work, we explore the impact of training data size within a single language, future work will focus on extending this framework to support multiple languages and exploring language-agnostic ASR metric approximation.
@@ 7 Limitations
Evaluation. While our evaluation setup is comprehensive, consisting of over 40 models and 14 datasets representing various acoustic and linguistic conditions such as natural noise, dialects, and accents—far surpassing previous works—we have not explored more nuanced conditions such as gender, non-native speech, and approximation across various age groups. Additionally, while the framework has demonstrated strong performance in approximating ASR metrics across multiple datasets, its generalization to highly diverse or extreme real-world conditions might still require further investigation.
Language. Furthermore, the evaluation is currently limited to a single language; expanding this framework to multiple languages or achieving language-agnostic ASR metric approximation remains an important direction for future work.
Compute. While, unlike previous works, our final approximator is a simple machine learning model that does not require GPUs to run, we do utilize a single GPU for multimodal embedding extraction, which could be performed on any consumer-grade GPU.
@@ A.3 Models
Nemo-ASR-Models Gulati et al. (2020); Variani et al. (2020); Rekesh et al. (2023); Noroozi et al. (2024); Tang et al. (2023); Harper et al. We included several NVIDIA’s NeMo advanced automatic speech recognition (ASR) models, including Canary, Parakeet (110M, 0.6B, and 1.1b), Conformer-CTC, and Fast-Conformer, as each is designed for specific use cases and optimized for performance. Canary-1B is a state-of-the-art multilingual, multitask model featuring a FastConformer encoder and Transformer decoder. The Parakeet family includes models with a FastConformer encoder paired with different decoders: CTC, RNN-T, or TDT. Conformer-CTC is a non-autoregressive model based on the Conformer architecture, combining self-attention and convolution for global and local feature extraction. It uses CTC loss and a linear decoder, supporting both sub-word (BPE) and character-level encodings. While Fast-Conformer is an optimized version of the Conformer architecture, offering significant speed improvements (2.4x faster) with minimal quality degradation. It uses 8x depthwise convolutional subsampling and reduced kernel sizes for efficiency.
@@ A.4 Experiments
The regression framework is a stacking ensemble comprising multiple base regressors and a final estimator. We perform basic hyperparameter tuning using RandomizedSearchCV with 5-fold cross-validation, with the objective to minimize mean absolute error (MAE). The search explores key hyperparameters such as n_estimators, max_depth, learning_rate, and min_samples_split, balancing model complexity and generalization. We provide hyperparameter and other details in 5. The model is trained on 14 datasets divided into two groups: bench (10 standard benchmark datasets) and in-the-wild (4 diverse, real-world datasets). A leave-one-out strategy is applied to the bench set, where the model is trained on 9 datasets and evaluated on the remaining one. All trained models are also evaluated on the in-the-wild set, which remains isolated during training to assess out-of-domain generalization.
@@
A.5 Results

### v2

@@ 1 Introduction
Automatic Speech Recognition (ASR) models have made significant advancements in recent years, achieving near-human performance on several standard evaluation benchmarks (Radford et al., 2022; Seamless Communication et al., 2023; Communication et al., 2023; Harper et al., 2024, inter alia). These models are typically evaluated using metrics like Word Error Rate (WER) and Character Error Rate (CER) Likhomanenko et al. (2020), which are essential for assessing model performance.
@@
  • We evaluate over 40 ASR models across 14 diverse evaluation setups, including both standard benchmarks and domain-specific, unseen conditions, followed by training regression models to approximate ASR metrics.
  • We compare our approach with the most recent work on approximating ASR metrics and show over a 50% reduction in absolute difference against the strong baseline.
@@ 2 Related Work
Automatic speech recognition (ASR) has witnessed significant advancements in recent years, primarily due to the scaling of both data and model size Radford et al. (2022); Communication et al. (2023). Transformer Vaswani et al. (2023) based models, in particular, have significantly contributed to these developments by effectively capturing long-range dependencies and contextual nuances in speech, achieving state-of-the-art (SOTA) performance across diverse benchmarks Kheddar et al. (2024); Dhanjal and Singh (2024); Zimerman and Wolf (2023). While traditional evaluation metrics like Word Error Rate (WER) and Character Error Rate (CER) are de-facto evaluation metrics in benchmarking ASR systems Lin et al. (2021); Park et al. (2024), scenarios where ground truth transcriptions are unavailable have caught interest in reference-free ASR evaluation methods Karbasi and Kolossa (2022); Wang et al. (2024); Kuhn et al. (2024).
Reference-free ASR evaluation methods aim to estimate ASR performance without requiring ground truth transcriptions Ospanov et al. (2024). Earlier approaches rely on heuristic features or metadata such as speaker demographics, background noise, and linguistic characteristics Litman et al. (2000); Yoon et al. (2010), limiting their applicability across varied contexts. However, recent advancements focus on deep learning-based frameworks, such as convolutional neural networks (CNNs) Elloumi et al. (2018) and contrastive learning methods Yuksel et al. (2023a), to predict ASR quality directly from encoded speech and text. For instance, methods like NoRefER Yuksel et al. (2023b) use Siamese architectures fine-tuned on ASR hypotheses, achieving high correlation with traditional metrics and improving WER by optimizing hypothesis ensembling Park et al. (2024).
@@ 3.4 Evaluation
We evaluate the regression model’s performance across four setups, including IID and OOD data and different model configurations. Specifically, we train our regression model on one ASR system (source) on one dataset and evaluate it on both IID and OOD data for the source and target models. We provide a detailed analysis of the distribution shift in Appendix A.5.
@@ 4.1 Datasets
To evaluate the robustness and generalizability of our ASR metrics approximation tool, we use datasets sourced from multiple distributions, divided into two types: Standard Benchmark and Wild Challenge datasets. We describe these datasets below and provide additional details in Appendix A.2, Table 5.
@@
In addition to the above datasets, we also run a small experiment to assess the cross-lingual transferability of the trained regression model. Specifically, we train the model on English data and evaluate it on both German and English data, and vice versa, using the English and German splits from LibriSpeech Panayotov et al. (2015) and Common Voice Ardila et al. (2020). In our ablation study, we compare the trained regression model with a proxy reference. To broaden this evaluation, we expand the in-the-wild dataset to include four additional datasets: a privately collected Medical ASR dataset with clinical conversations; standard data with eight synthetic perturbations (white noise, time stretch, pitch shift, cross-lingual noise, reverberation, pub noise, echo, and distortion); noisy home recordings BERSt Tuttösí et al. (2025); and CHiME-6(noisy subset) Watanabe et al. (2020).
@@ 4.2 Models
Encoder-Decoder Models. We include multiple encoder-decoder families of models capable of performing ASR tasks in a zero-shot setting. More specifically, we include whisper Radford et al. (2023) and distil-whisper Gandhi et al. (2023) models that perform really well across diverse testing settings. We also include seamless  Communication et al. (2023); Seamless Communication et al. (2023); Barrault et al. (2025), SpeechT5 Ao et al. (2022) which are unified encoder-decoder frameworks for tasks such as ASR, speech synthesis, translation, and voice conversion. MMS Pratap et al. (2023) supports hundreds of languages and excels in resource-constrained scenarios. Moonshine (2) Jeffries et al. (2024), a lightweight and efficient model, is designed for edge deployments with strong performance. Additionally, we include speech language models like SpeechLLM Rajaa and Tushar , which combine speech embeddings with language models to predict metadata such as speaker attributes, emotions, and accents, offering robust multimodal capabilities.
NeMo-ASR Models. We use multiple models from the NeMo-ASR Gulati et al. (2020); Variani et al. (2020); Noroozi et al. (2024); Tang et al. (2023); Harper et al. (2024) toolkit by NVIDIA. We include models such as Canary and Parakeet, which use highly efficient speech encoders like Fast-Conformer Rekesh et al. (2023). In addition to that, we use models based on various encoders and decoders (CTC, RNN-T, TDT, Conformer-CTC Guo et al. (2021). In our work, we evaluate 11 models from the NeMo-ASR toolkit.
Encoder-Only. We include self-supervised encoder-only models and their derivatives. Specifically, we use Wav2Vec2 Schneider et al. (2019); Baevski et al. (2020), HuBERT Hsu et al. (2021), and Data2Vec Baevski et al. (2022).
@@ 4.3 Experimental Setup
For multimodal embeddings, we use SONAR Duquenne et al. (2023), a 1024-dimensional sentence-level multilingual model. Specifically, we utilize text_sonar_basic_encoder for text encoding and speech_sonar_basic_encoder for speech encoding.
@@ 5 Results
We conduct experiments using two dataset categories: standard benchmarks and in-the-wild, as described in Section 4.1. For each ASR model, a leave-one-out strategy is used, training the regression model on 9 benchmark datasets and testing it on the remaining benchmark dataset and all four in-the-wild datasets to ensure comprehensive evaluation on out-of-domain data. Additionally, in-domain testing is included in ablation studies, as detailed in Section 5.4. The regression model is trained to predict absolute error counts (word and character levels), which are normalized by the reference length to compute approximate error rates ($aWER$ and $aCER$). We also train regression models to directly predict WER and CER. We provide results for fine-grained metrics in Table 12.
@@ 5.1 Evaluation on In-the-Wild Datasets
The wild datasets provide a realistic testbed for evaluating the regression model’s ability to approximate error rates under real-world conditions. As shown in Table 1, high-performing models, like canary-1b, demonstrate strong agreement between predicted and actual error rates. For example, on VP_Accented, canary-1b achieves mean absolute difference of $1.1\%$. On Primock57, the model shows robustness with a WER of $16.2\%$ and an $aWER$ of $13.4\%$, highlighting its effective generalization across diverse and domain-specific contexts.
For models like data2vec-audio-large-960h our approximation is pretty close to actual error rates with difference consistently under $2\%$ on various datasets. For example, on LibriSpeech-test-noise, the model’s actual WER is $7.2\%$ while the approximated $aWER$ is $8.6\%$. Even on acoustically complex datasets like ATCOsim, where the WER is $44.0\%$ and the $aWER$ is $51.1\%$, the model exhibits a reasonable alignment between approximated and actual error rates.
@@ 5.2 Evaluation on Benchmark Datasets
We summarize results on 10 standard benchmark datasets in Appendix A.6 Tables 13 and 14. Each table reports actual WER/CER alongside the approximated WER/CER (denoted by aWER/aCER).
@@
Figure 2: Actual and approximated WER for four models across standard benchmark.
@@
5.3 Multilingual Evaluation
We train the regression model on English and evaluate it on English and German, and vice versa. We do this experiment with two models as source and proxy, namely seamless-m4t-v2-large and whisper-large-v3. We report the mean absolute difference between approximated and actual word error counts in Table 2.
[…]
Table 2: Cross-lingual mean absolute difference ($\downarrow$) between predicted and actual word error counts. Lower values mean better approximation.
We find that our framework demonstrates strong cross-lingual generalization: when trained on English data, the regression model maintains low absolute differences when evaluated on German datasets, and vice versa. This consistency across languages and datasets, using both seamless-m4t-v2-large and whisper-large-v3 as source-proxy pairs, underscores the robustness and language-agnostic nature of our approach. These results validate that our method can be effectively applied in multilingual settings without the need for language-specific adaptations.
@@ 5.4 Ablation
We conduct comprehensive ablation experiments to evaluate the robustness of the approximation model and the contributions of its individual components. Using the evaluation setup outlined in Section 3.4, we select data2vec-audio-base-960h as the source model ($S$) and wav2vec2-base-960h as the target model ($T$). The results are summarized in Table 5.4, where IID results correspond to Case-I 3.4, and $D$, $M$, and $D+M$ under OOD represent Case II 3.4, Case-III 3.4, and Case-IV 3.4, respectively. The reference model’s $r$ value represents the average WER across all datasets. We include reference models with varying $r$ values, such as whisper-large-v3 ($r=17.8$), whisper-medium.en ($r=20.1$), whisper-tiny ($r=33.4$), and mms-1b-fl102 ($r=51.0$).
@@ D M D + M
The quality of references, quantified by the $r$-value, also plays a critical role. For example, in IID conditions, the MAE increases from $1.31$ for $r=17.8$ to $2.03$ for $r=51.0$. A similar trend is observed in OOD $D+M$, where the MAE rises from $1.40$ ($r=17.8$) to $2.09$ ($r=51.0$). The absence of similarity (w/o S) combined with low-quality proxies further degrades performance, underscoring the importance of both high-quality references and similarity measures. We provide character-level error count approximation in Appendix A.6 Table 10.
@@
Direct Comparison with Proxy Reference.
We evaluate our regression model and standard reference-free baseline that computes the word error rate between the target and a proxy hypothesis. We report the mean absolute difference and compare the regression model (Ours) with direct calculation with proxy (w Proxy) in Table 4.
Our results show that the regression model consistently reduces the absolute difference compared to direct evaluation with the proxy across seven of the eight datasets, with the only exception being MedicalASR. We also find that the regression model trained using only similarity features remains competitive. These findings demonstrate that our approach generalizes well across domains and provides more reliable reference-free ASR quality estimates than simply computing the error against a proxy reference.
Dataset PR (r=33.4) PR (r=51.0) w/o Proxy
w Proxy Ours w Proxy Ours Ours
Primock57 2.25 1.54 4.89 2.63 3.82
ATCOSim 5.35 2.11 3.56 2.02 2.64
VP_Accented 2.99 2.09 5.39 2.83 4.00
LS_Noise 2.10 1.58 3.08 1.42 2.35
BERSt 2.37 1.25 5.09 2.47 3.20
Perturbed 1.84 1.34 3.38 1.75 2.81
Medical 0.76 0.95 2.35 1.33 2.56
CHiME6-Noisy 2.32 1.43 3.29 1.93 2.74
Average 2.50 1.54 3.88 2.05 3.02
Table 4: Mean absolute difference ($\downarrow$) between the predicted and actual word error counts for our regression models and the baseline direct comparison with proxy reference (w Proxy).
@@ 6 Conclusion
We present a framework for approximating ASR metrics, demonstrating its effectiveness in generalizing to unseen, in-the-wild, and challenging conditions. Our results show that the model performs well with absolute error counts, consistently outperforming strong baseline, with error rates remaining relatively low. We show that our proposed method achieves consistent performance across 40 ASR models and 14 evaluation setups, including both standard benchmarks and domain-specific conditions. The trained regression model can be efficiently used to approximate ASR metrics, particularly in data-constrained environments, such as critical domains with limited labeled data. In summary, our work bridges the gap between theoretical advancements and real-world applications, paving the way for more reliable and scalable ASR systems. While in this work, we evaluate monolingual and cross-lingual generalization, future work will focus on extending this framework to support a multilingual setting and exploring language-agnostic ASR metric approximation.
@@ 7 Limitations
Evaluation. While our evaluation setup is comprehensive, consisting of over 40 models and 14 datasets representing various acoustic and linguistic conditions such as natural noise, dialects, and accents—far surpassing previous works—we have not explored more nuanced conditions such as gender, non-native speech, and approximation across various age groups. Additionally, while the framework has shown strong performance in approximating ASR metrics across multiple datasets, its generalization to highly diverse or extreme real-world conditions might still require further investigation.
Language. Additionally, the current evaluation focuses solely on monolingual and bilingual settings. Extending this framework to include multiple languages and rigorously testing it across diverse linguistic contexts represents a critical direction for future research.
Compute. Unlike previous works, our final approximator is a simple regression model that does not require GPUs to run, we do utilize a single GPU for multimodal embedding extraction, which could be performed on any consumer-grade GPU.
@@ A.3 Models
Nemo-ASR-Models Gulati et al. (2020); Variani et al. (2020); Rekesh et al. (2023); Noroozi et al. (2024); Tang et al. (2023); Harper et al. (2024) We included several NVIDIA’s NeMo advanced automatic speech recognition (ASR) models, including Canary, Parakeet (110M, 0.6B, and 1.1b), Conformer-CTC, and Fast-Conformer, as each is designed for specific use cases and optimized for performance. Canary-1B is a state-of-the-art multilingual, multitask model featuring a FastConformer encoder and Transformer decoder. The Parakeet family includes models with a FastConformer encoder paired with different decoders: CTC, RNN-T, or TDT. Conformer-CTC is a non-autoregressive model based on the Conformer architecture, combining self-attention and convolution for global and local feature extraction. It uses CTC loss and a linear decoder, supporting both sub-word (BPE) and character-level encodings. While Fast-Conformer is an optimized version of the Conformer architecture, offering significant speed improvements (2.4x faster) with minimal quality degradation. It uses 8x depthwise convolutional subsampling and reduced kernel sizes for efficiency.
@@ A.4 Experiments
The regression framework is a stacking ensemble comprising multiple base regressors and a final estimator. We perform basic hyperparameter tuning using RandomizedSearchCV with 5-fold cross-validation, with the objective to minimize mean absolute error (MAE). The search explores key hyperparameters such as n_estimators, max_depth, learning_rate, and min_samples_split, balancing model complexity and generalization. We provide hyperparameter and other details in 7. The model is trained on 14 datasets divided into two groups: bench (10 standard benchmark datasets) and in-the-wild (4 diverse, real-world datasets). A leave-one-out strategy is applied to the bench set, where the model is trained on 9 datasets and evaluated on the remaining one. All trained models are also evaluated on the in-the-wild set, which remains isolated during training to assess out-of-domain generalization.
@@
A.5 Domain Divergence and Phonetic Diversity Analysis
To quantify how our in-the-wild evaluation sets differ from the LibriSpeech-clean corpus that was used to fine-tune both data2vec-audio-base-960h (source) and wav2vec2-base-960h (target), we measure acoustic domain divergence with Central Moment Discrepancy (CMD) computed over SONAR speech embeddings and phonetic diversity with Total Vocabulary Overlap (TVO) calculated on the corresponding transcripts. We report these numbers in Tables 8 and 9.
Dataset CMD Interpretation
LibriSpeech (Noise) 0.062 Low divergence
Primock57 0.329 Significant divergence
VoxPopuli (Accented) 0.504 High divergence
ATCOSIM 0.538 High divergence
Table 8: Acoustic domain divergence between LibriSpeech-clean and each evaluation set, measured with Central Moment Discrepancy (CMD).
Dataset TVO (%) Interpretation
LibriSpeech (Noise) 36.9 High overlap
VoxPopuli (Accented) 20.3 Low overlap
Primock57 14.3 Low overlap
ATCOSIM 3.8 Very low overlap
Table 9: Total Vocabulary Overlap (TVO) between LibriSpeech-clean and each evaluation set.
CMD values confirm that LibriSpeech (Noise) remains acoustically close to the source domain because only artificial background sounds are added, whereas Primock57, VoxPopuli (Accented), and ATCOSIM exhibit substantial acoustic shifts. TVO scores tell a complementary story in the lexical space: LibriSpeech (Noise) preserves more than one-third of the vocabulary, while Primock57, VoxPopuli (Accented), and especially ATCOSIM share far less with LibriSpeech-clean. Together, these metrics justify the in-the-wild designation of our evaluation corpora and highlight the importance of robust models that generalize beyond clean, studio-quality speech.
A.6 Results